at constant pressure,what would be the percentage decrease in the density of an ideal gas for a 10% increase in the temperature?

Grade:11

1 Answers

Arun
3 years ago
Pressure is contant (given), M.W (Molecular Weight) is obviously constant
 
Now Temperature increase is 10%
 
\rightarrow P(M.W)=\sigma (\frac{x}{100})R(\frac{110}{100}T)
\rightarrow (\frac{x}{100})(\frac{110}{100})=1\ (No\ Change)
\rightarrow x=90.9\%
 
Therefore there is a decrease of 9.1% (9 approx)